# How Do Companies Manage the Environmental Sustainability of AI? An Interview Study About Green AI Efforts and Regulations

### v1

@@ V Discussion
Regarding RQ1, the first theme revealed that the primary driver for adopting AI was efficiency. Other mentioned drivers were security, personal passion, and the competitive market surrounding AI. This is in line with the findings of Radhakrishnan and Chattopadhyay [40], where the facilitating factors of AI adoption were performance efficacy and efficiency, intrinsic motivation, and customer needs. Once AI is adopted in the workplace, it is used for various tasks, such as data processing or as a chatbot. AI has a significant contribution to many people’s work lives, so it is crucial to address its downsides to maximize its benefits. During the adoption of AI, there was an awareness of the negative environmental impact of these AI tools, particularly of its carbon footprint and energy consumption. It was only noted once that AI models also contribute to global water shortages [26], which supports previous findings on the narrow focus on carbon reduction in climate change discussions [50].
@@ VI Limitations
Finally, the popularity or desirability of AI may have influenced the participants’ responses, particularly concerning sustainability awareness and efforts. Some interviewees, especially those in leadership positions, may have framed their company’s approach to AI and environmental sustainability in a more favorable light, consciously or unconsciously aligning their answers with perceived best practices. This concurs with research indicating that professionals often tailor their responses to align with expected corporate narratives [51]. Future studies could mitigate this limitation by incorporating observational data or case studies alongside self-reported interviews. Nonetheless, the small number of applied Green AI practices was explicitly recognized by the participants, suggesting that they answered fairly truthfully.

### v2

@@ V Discussion
Regarding RQ1, the first theme revealed that the primary driver for adopting AI was efficiency. Other mentioned drivers were security, personal passion, and the competitive market surrounding AI. This is in line with the findings of radhakrishnan2020determinants, where the facilitating factors of AI adoption were performance efficacy and efficiency, intrinsic motivation, and customer needs. Once AI is adopted in the workplace, it is used for various tasks, such as data processing or as a chatbot. AI has a significant contribution to many people’s work lives, so it is crucial to address its downsides to maximize its benefits. During the adoption of AI, there was an awareness of the negative environmental impact of these AI tools, particularly of its carbon footprint and energy consumption. It was only noted once that AI models also contribute to global water shortages [george2023environmental], which supports previous findings on the narrow focus on carbon reduction in climate change discussions [savasta2014dangers].
@@ VI Limitations
Finally, the popularity or desirability of AI may have influenced the participants’ responses, particularly concerning sustainability awareness and efforts. Some interviewees, especially those in leadership positions, may have framed their company’s approach to AI and environmental sustainability in a more favorable light, consciously or unconsciously aligning their answers with perceived best practices. This concurs with research indicating that professionals often tailor their responses to align with expected corporate narratives [paulhus2002use]. Future studies could mitigate this limitation by incorporating observational data or case studies alongside self-reported interviews. Nonetheless, the small number of applied Green AI practices was explicitly recognized by the participants, suggesting that they answered fairly truthfully.